A person is driving a car with speed of vo. The person brakes to stop over time t and distance d. The magnitude of the person acceleration is a. If the person doubles her initial velocity , how much distance does she travels as she stops
1
Expert's answer
2020-09-21T08:30:46-0400

The distance traveled by a car to the rest

"d=\frac{v_0^2}{2a}"

If the person doubles her initial velocity then she's distance traveled to the stop increases by factor four:

"d'=\frac{(2v_0)^2}{2a}=4\frac{v_0^2}{2a}=4d"